OpenStudy (anonymous):

Solve for "y" 2x+y=5

OpenStudy (shamil98):

solve for y then. atleast show your attempt.

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

Y=5-2x

OpenStudy (anonymous):

how did you get that answers

OpenStudy (anonymous):

y=5-2x

OpenStudy (anonymous):

you isolate y so in the origional problem 2x+y=5 you subtract 2x from both sides and then you get y=5-2x

OpenStudy (anonymous):

thanks you....

OpenStudy (anonymous):

youre welcome hope it was helpful